You are currently working on a project that had an original budget of $300,000 and an original schedule of 27 weeks. You currently have spent $100,000 and are over budget by $15,000. Your boss comes into your office and wants to discuss your project. Specifically, she wants to know how much more money it is going to take to complete the project. Assuming you have completed $78,000 of the deliverables. What do you tell her?
Answer #1

Given: Budget at completion BAC = $300,000

Actual cost = AC = $100,000

Overbudget = $15,000

Planned Value PV = AC - Overbudget = 100,000 - 15,000 = 85,000

Work completed = Earned Value = EV = $78,000

Cost performance index = CPI = EV / AC = 78000 / 100000 = 0.78

Amount required more = Estimate to Complete = = = $284,615.38

Hence, We tell her that $284,615.38 are required more to complete the project.

____________________________________________________________________________________________

In case of any doubt, please ask through comment section before Upvote / downvote.

Kindly rate the answer as it will be encouraging for me to keep answering further questions!!!